OpenStudy (mimi_x3):

Only for Kaffeine, easy one first. The cost to run a vehicle is a + v^2/a dollars per hour where v is the average speed of the vehicle in km/h. Show thatt the cost for a trip of Dkm is lowest if the vehicle travels at v=a km/h

OpenStudy (katrinakaif):

Do you want the work to? O_o

OpenStudy (mimi_x3):

nope, wait its a show question. Just show the working if you're bothered.

OpenStudy (katrinakaif):

Okay sorry, the stupid thing said server lost connection C = a + v2/a => dc/dv = 0 + 2v/a => dc/dv = 2v/a => 2v /a = 0 2v = 0(a) v = 0 => d2c / dv2 = d/v(2v/a)= 2/a (Postive value so cost is lowest for all values of v ) ??

OpenStudy (mimi_x3):

Sorry, its wrong.

OpenStudy (katrinakaif):

K give me a moment.

OpenStudy (mimi_x3):

You can ask for help if you want, i don't mind.

OpenStudy (katrinakaif):

Omfg Mimi.

OpenStudy (mimi_x3):

lols, what's wrong ? Archaic Mathematician can help :P

OpenStudy (anonymous):

This is basic calc (calc 1), i wanna see her do this.

OpenStudy (mimi_x3):

Isn't calc 1 ? college ? >_<

OpenStudy (anonymous):

she can do it, i am sure. She solved harder ones before

OpenStudy (mimi_x3):

ok, i'll give a little hint, it involved d s t

OpenStudy (mimi_x3):

involves*

OpenStudy (katrinakaif):

c(a) = 2v?? don't tell me I am wrong again. -_-

OpenStudy (mimi_x3):

lols, yeah sorry. Want me to type up the solution ?

OpenStudy (katrinakaif):

Lol Give me the answer and I ll figure the solution

OpenStudy (anonymous):

no way

OpenStudy (mimi_x3):

It's a show.

OpenStudy (mimi_x3):

No way ?

OpenStudy (katrinakaif):

Lagrange, stay out of it, and go away.

OpenStudy (katrinakaif):

Mimi, type.

OpenStudy (mimi_x3):

Archaic mathematician can help you.

OpenStudy (katrinakaif):

I ll figure it out myself.

OpenStudy (anonymous):

hard headed ^

OpenStudy (mimi_x3):

lols, and stubborn -_-

OpenStudy (katrinakaif):

Shut up featheres.

OpenStudy (anonymous):

lets get some soap for that mouth

OpenStudy (mimi_x3):

Please refrain from swearing, don't get angry man.

OpenStudy (katrinakaif):

Lol You have no control over my fingers, so no use in stating stop :P Please feel free to hit the report abuse button =) It will be a donation

OpenStudy (mimi_x3):

Hm...dw about the problem, I'll type up the solution if you want.

OpenStudy (anonymous):

yeah, put her out her misery

OpenStudy (mimi_x3):

lols, yeah i feel bad.

OpenStudy (katrinakaif):

Didn't I say I ll figure it out on my own?

OpenStudy (mimi_x3):

It'll take a while, it took me a while.

OpenStudy (katrinakaif):

Doesn't matter. I have a brain gifted to me for it.

OpenStudy (mimi_x3):

Okay, do it as you wish, don't need to give me the solution straight away.

OpenStudy (anonymous):

yeah, let her use "her" brain to find the solution

OpenStudy (katrinakaif):

Lagrange, how many effing times Have I told you to stay away from me? Gosh damn.

OpenStudy (mimi_x3):

lols, dont be mean :P

OpenStudy (mimi_x3):

Kaffeine is pissed off

OpenStudy (anonymous):

woah, relax buddy

OpenStudy (anonymous):

take a chill pill

OpenStudy (katrinakaif):

Get a life dude, other than critizing me each time I solve a calc problem.

OpenStudy (katrinakaif):

You crossed the limits a while ago, and I told you to STAY AWAY

OpenStudy (anonymous):

nobody is criticizing you, i just find it weird that you can solve problems i give you, that are straight from calc 2, but you cant solve this one

OpenStudy (mimi_x3):

You are making it sound so serious. Don't worry about it. I'll take up the solution.

OpenStudy (katrinakaif):

Lol Does a person have be good at calc 1 to get into calc bc? Your derivative questions are easy as hell.

OpenStudy (mimi_x3):

Time to travel D km T = D / S = D / V (hours) Cost to travel D km C = T * ( a + v^2/a ) = D / v ( a + v^2 / a ) = D ( a/v + v/a) dC / dt = D( -a/v^2 + 1/a) = 0 -a/v^2 + 1/a = 0 1/a = a/v^2 v^2 = a^2 -> a km/h check for minimum: (using second derivative) d^2C/dt^2 = D(2a/v^3 + 0) = D * 2a/a^3 = 2D/a^2 > 0 yes, the cost at v=a is the lowest. Therefore v=a km/h for lowest cost.

OpenStudy (katrinakaif):

Damn. forgot to plug it into the equation

OpenStudy (mimi_x3):

Yeah and you forgot about the distance.

OpenStudy (katrinakaif):

Pellet >.>

OpenStudy (katrinakaif):

I used the distance to get my 2 values in the first one..Forgot all about it

OpenStudy (mimi_x3):

Yeah, when i did it I forgot about the other one as well xD

OpenStudy (anonymous):

it was a nice try

OpenStudy (mimi_x3):

Want an easier one kaffeine ?

hero (hero):

This is all hilarious @Mimi_x3 @LagrangeSon678 @KatrinaKaif